# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v2

@@ 9.5.3 Learning-Based Generation of a Sequence of Contact Points and Forces/Torques
Hand motion capture can be categorized into camera-based, IMU-based, and electromagnet-based methods. Using camera-based methods, human motion is estimated through pose estimation and retargeted to a multi-fingered end-effector. However, the accuracy of pose estimation needs improvement and occlusions can cause information loss. IMU-based methods are typically used to estimate wrist pose but miss fine-grained finger motion. Additionally, IMU used for pose estimation are prone to drifting over time. Electromagnet-based methods primarily use electromagnetic field gloves to estimate finger motion, but they lack information about wrist pose. At the same time, IMU-based methods and electromagnet-based methods both lack the vision data needed for policy training. DEXCAP (Wang et al. (2024)) designs a portable hand motion capture system. It uses Rokoko gloves to capture finger motion. A Realsense T265 camera is added to the wrist of the gloves to monitor wrist pose. Additionally, a Realsense L515 LiDAR camera is mounted on the chest to sense the environment. However, the glove’s design lacks tactile feedback. Therefore, to collect data for a multi-fingered end-effector, hardware improvements are necessary.
Two main learning-based methods for dexterous manipulation are imitation learning (Ze et al. (2024)) and reinforcement learning (Ma et al. (2023b)). Imitation learning can use a visual encoder (in Sec. 6) for visuo-motor control. Diffusion policy (Chi et al. (2023)) adapts the concept of diffusion to visuo-motor control. It addresses challenges in visuo-motor control such as action multimodality, sequential correlation to accommodate high-dimensional action sequences. However, it remains uncertain whether diffusion policy is suitable for one-model to multiple tasks approach. It can also use an existing VLM for fine-tuning (in Sec. 7). Fine-tuning with a VLM allows a skill to work in an open world. This often performs better on unseen objects compared to visuo-motor control (Brohan et al. (2023)). Reinforcement learning offers exploration capability, which address suboptimal issues. This advantage distinguishes it from imitation learning. However, reinforcement learning is primarily trained in simulation. It still has limitations in addressing the sim-to-real challenge of complex tasks, such as pen-spinning. In Sec. 7, the use of foundation models to assist reinforcement learning is introduced. FAC (Ye et al. (2023a)) offers potential for training reinforcement learning in real-world environment, but it still lacks consideration of environment resets (Gupta et al. (2021)) and safety. Therefore, using foundation models to assist reinforcement learning in real-world training requires further exploration.
@@ 9.5.4 Whole-body Control
For whole-body control, the focus is on low-level control issues. A straightforward idea is to expand the action space of the policy module’s robot foundation models to include all joints of the robot. However, as the output dimensions increase, end-to-end training methods are more likely to diverge. Therefore, most current models output cartesian space poses and force/torques. These outputs are then optimized and converted into position or torque for each joint through a post-processing module (Haviland and Corke (2021)). To address end-to-end whole-body control issues, principal research is needed to facilitate network training and deployment.
@@ 9.6 How to establish a benchmark?
Table. 1 lists the benchmarks used in current RFMs, highlighting a lack of standardization. This inconsistency hinders the development of RFMs for three main reasons. Firstly, current RFMs are tied to the specific parameters of each robot, such as the choice of sensors, camera pose, and the robot’s degrees of freedom. These factors prevent RFMs from being easily transferred across different robots. Secondly, testing the generalization and success rate of general manipulation capability requires a wide range of scenes and tasks. Thirdly, there is no standardized metric for assessing general manipulation capability.
@@ 10 Conclusion
Then, we discuss the following points: (1) the logic and implementation strategies of the designed framework, (2) how to design the modules of post-conditions detection and post-hoc correction, due to limited research on foundation models for these areas, (3) the learning capability required for general manipulation, (4) the strengths and weaknesses of current RFMs, (5) how to make the general manipulation with the dexterous capability, and (6) how to establish a foundation models for manipulation benchmark.
Additionally, the proposed framework has certain limitations: (1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates and safety for specific tasks. Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework is limited to static environments, rigid objects, short-horizon tasks, improving old skills, and low precision requirements for contact points and forces/torques. Further research is needed to advance from L0 to L5.

### v3

@@ 9.9.3 Learning-Based Generation of a Sequence of Contact Points and Forces/Torques
Hand motion capture can be categorized into camera-based, IMU-based, and electromagnet-based methods. Using camera-based methods, human motion is estimated through pose estimation and retargeted to a multi-fingered end-effector. However, the accuracy of pose estimation needs improvement and occlusions can cause information loss. IMU-based methods are typically used to estimate wrist pose but miss fine-grained finger motion. Additionally, IMU used for pose estimation are prone to drifting over time. Electromagnet-based methods primarily use electromagnetic field gloves to estimate finger motion, but they lack information about wrist pose. At the same time, IMU-based methods and electromagnet-based methods both lack the vision data needed for policy training. DEXCAP (Wang et al. (2024b)) designs a portable hand motion capture system. It uses Rokoko gloves to capture finger motion. A Realsense T265 camera is added to the wrist of the gloves to monitor wrist pose. Additionally, a Realsense L515 LiDAR camera is mounted on the chest to sense the environment. However, the glove’s design lacks tactile feedback. Therefore, to collect data for a multi-fingered end-effector, hardware improvements are necessary.
Two main learning-based methods for dexterous manipulation are imitation learning (Ze et al. (2024a)) and reinforcement learning (Ma et al. (2023b)). Imitation learning can use a visual encoder (in Sec. 6) for visuo-motor control. Diffusion policy (Chi et al. (2023)) adapts the concept of diffusion to visuo-motor control. It addresses challenges in visuo-motor control such as action multimodality, sequential correlation to accommodate high-dimensional action sequences. However, it remains uncertain whether diffusion policy is suitable for one-model to multiple tasks approach. It can also use an existing VLM for fine-tuning (in Sec. 7). Fine-tuning with a VLM allows a skill to work in an open world. This often performs better on unseen objects compared to visuo-motor control (Brohan et al. (2023)). Reinforcement learning offers exploration capability, which address suboptimal issues. This advantage distinguishes it from imitation learning. However, reinforcement learning is primarily trained in simulation. It still has limitations in addressing the sim-to-real challenge of complex tasks, such as pen-spinning. In Sec. 7, the use of foundation models to assist reinforcement learning is introduced. FAC (Ye et al. (2023a)) offers potential for training reinforcement learning in real-world environment, but it still lacks consideration of environment resets (Gupta et al. (2021)) and safety. Therefore, using foundation models to assist reinforcement learning in real-world training requires further exploration.
@@ 9.9.4 Whole-body Control
For whole-body control, the focus is on low-level control issues. A straightforward idea is to expand the action space of the policy model to include all joints of the robot. However, as the output dimensions increase, end-to-end training methods are more likely to diverge. Therefore, most current models output cartesian space poses and force/torques. These outputs are then optimized and converted into position or torque for each joint through a post-processing module (Haviland and Corke (2021)). To address end-to-end whole-body control issues, principal research is needed to facilitate network training and deployment.
@@ 9.10 How to establish a benchmark?
Tab. 1 lists the benchmarks used in current RFMs and we list some representative benchmarks in Tab. 3, highlighting a lack of standardization. This inconsistency hinders the development of RFMs for three main reasons. Firstly, current RFMs are tied to the specific parameters of each robot, such as the choice of sensors, camera pose, and the robot’s degrees of freedom. These factors prevent RFMs from being easily transferred across different robots. Secondly, testing the generalization and success rate of general manipulation capability requires a wide range of scenes and tasks. Thirdly, there is no standardized metric for assessing general manipulation capability.
@@ 10 Conclusion
Then, we discuss the following points: (1) the logic and implementation strategies of the designed framework, (2) how to design the modules of post-conditions detection and post-hoc correction, due to limited research on foundation models for these areas, (3) the learning capability required for general manipulation, (4) the strengths and weaknesses of current RFMs, (5) how to use internet-scale video data for RFMs, (6) how to design low-cost hardware for collecting large-scale real-world manipulation data, (7) how to collect large-scale diversity and realism manipulation data in simulation, (8) comparative analysis of 2D and 3D-based methods, (9) how to make the general manipulation with the dexterous capability, and (10) how to establish a foundation models for manipulation benchmark.
Additionally, the proposed framework has certain limitations: (1) The framework is designed with a sequential structure, which contrasts with the parallel execution in human operation. (2) Both the proposed framework and the surveyed literature are based on learning-based approaches. While model-based methods may not generalize as well, they tend to significantly outperform learning-based methods in terms of success rates, precision and safety for specific tasks (Pang et al. (2023)). Therefore, investigating the integration of learning-based and model-based approaches remains an important research. (3) The framework proposed in this paper is based on the development of learning-based methods and the definition of general manipulation. The framework of brain-like cognitive research should also be explored.